what is the concentration of each ion in the solution when 47.3 g of aluminum nitrate (Al(NO3)3) is dissolved in 1400 mL of water


1
Expert's answer
2022-05-30T14:10:03-0400

1. We find the amount of substance (mol) of Al(NO3)3 . To do this, we divide the mass of aluminum nitrate by its molecular mass (213gr/mol):

n = 47,3 / 213 = 0,222 mol

2. Al(NO3)3 => Al3+ + 3 NO3

1mol — 1mol — 3mol

0,222mol — X mol — Y mol

X = 0,222 * 1 / 1 = 0,222 mol Al3+

Y= 0,222 * 3 / 1 = 0,666 mol NO3

3. We find the concentration of molarity of each ion. To do this, we divide the mole of the ion (n) by the volume (litres) of solution:

CM (Al3+) = 0,222 mol / 1,4 l = 0,159 mol/l

CM (NO3) = 0,666 mol / 1,4 l = 0,476 mol/l

ANSWER: Al3+ = 0,159 mol/l

NO3 = 0,476 mol/l